# CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation

### v2

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model. Various experiments and ablations are done on four datasets including both the C++ and python languages to validate the hard meta-graph prompt, the soft prompting technique, and the effectiveness of the objectives for pretrained GNN expert. CodeGRAG improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. The implementation is available at https://anonymous.4open.science/r/Code-5970/ .
@@ 2.2 Graphical Knowledge Base Preparation
Abstract Syntax Tree (AST). An abstract syntax tree (AST) is a tree data structure that represents the abstract syntactic structure of source code. An AST is constructed by a parser, which reads the source code and creates a tree of nodes. Each node in the tree represents a syntactic construct in the source code, such as a statement, an expression, or a declaration. ASTs are used in a variety of compiler construction and program analysis tasks, including: parsing, type checking, optimization, and program analysis. ASTs have good compactness and can represent the structure of the source code in a clear and concise way.
@@
An illustration of the composed graphical view is given in Figure 3. After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
@@ 2.3 Knowledge Querying
Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base. The process can be formulated as:
$q\xleftarrow{}\text{QueryExtractor}(p),$ (3)
[…]
where $q$ denotes the query content, $p$ denotes the target problem, and $i$ is the returned index of the Top-1 relevant content stored in the constructed knowledge base.
The main problems of the retrieval lie in: 1) how to design the informative query content and 2) how to align the different modalities.
2.3.1 Query Extractor
Since the styles of different code problems can diversify, the query content of the retrieval process matters. We consider the following contents: 1) Problem description, which describes the task to be completed by the target function code. Potential ambiguity and style diversity may exist among different problems set, which lead to a decrease in retrieval accuracy. 2) Function declaration, which gives the function name and the input variables.
Before knowledge querying, we first extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
2.3.2 Retriever
The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
@@ 2.4.1 Hard Meta-Graph Prompt
The original graphical view of a code block could contain hundreds of nodes and edges. A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph template to abstract the information of the graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun & Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 5, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
Figure 4: Hard meta-graph prompt.
Figure 5: Soft prompting.
@@ 2.4.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components.
To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
@@
Finetuning with the Expert Soft Signal. To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5.
Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5.
@@ 3 Experiments
3.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website. For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1.
3.2 Main Results
The main results are summarized in Table 1 and Table 2.
@@ 3.2 Main Results
From the results, we can draw the following conclusions:
@@ 4 Related Work
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang & Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model. As retrieval augmented generation (RAG) shows its significance in promoting the quality of LLMs generation, works in code RAG start to accumulate. Nashid et al. (2023); Lu et al. (2022) utilize the code blocks as the retrieved knowledge to inform the LLMs with similar code blocks for enhancement. Zhou et al. (2022) uses the programming related document to serve as the retrieval content, injecting auxiliary external programming knowledge into the LLMs generation.
@@ 5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax and their multi-lingual code generation capabilities. This paper introduces the Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance LLMs for single round and cross-lingual code generation. CodeGRAG extracts and summarizes data flow and control flow information from codes, effectively bridging the gap between programming language and natural language. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency.
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
@@ Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
You may include other additional sections here.

### v3

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
Kounianhua Du1, Jizheng Chen1, Renting Rui1, Huacan Chai1, Lingyue Fu1, Wei Xia2, Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2 Huawei Noah’s Ark Lab Shanghai, China {kounianhuadu, wnzhang}@sjtu.edu.cn
@@ 2.2 Graphical Knowledge Base Preparation
Abstract Syntax Tree (AST). An abstract syntax tree (AST) is a tree data structure that represents the abstract syntactic structure of source code. An AST is constructed by a parser, which reads the source code and creates a tree of nodes. Each node in the tree represents a syntactic construct in the source code, such as a statement, an expression, or a declaration. ASTs have good compactness and can represent the structure of the source code in a clear and concise way.
@@
An illustration of the composed graphical view is in Figure 3. After obtaining the composed syntax graphs, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
@@ 2.3 Knowledge Querying
Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base.
We extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained. The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
@@ 2.4.1 Hard Meta-Graph Prompt
The original graphical view of a code block could contain hundreds of nodes and edges. A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph template to abstract the information of the graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 5 in the Appendix, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
@@ 2.4.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components. To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
@@
Finetuning with the Expert Soft Signal. To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5 in the Appendix.
Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5 in the Appendix.
@@ 3 Experiments
3.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website. For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1. More details of the retrieval pool and the finetuning setting can be found in Section A in the Appendix.
3.2 Main Results
The main results are summarized in Table 1 and Table 2. From the results, we can draw the following conclusions.
@@ 4 Related Work
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model. As retrieval augmented generation (RAG) shows its significance in promoting the quality of LLMs generation, works in code RAG start to accumulate. Nashid et al. (2023); Lu et al. (2022) utilize the code blocks as the retrieved knowledge to inform the LLMs with similar code blocks for enhancement. Zhou et al. (2022) uses the programming related document to serve as the retrieval content, injecting auxiliary external programming knowledge into the LLMs generation.
@@ 5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax. General large language models trained mainly on sequential-based natural language cannot well understand the structural-based programming language, e.g., the branching and jumping in codes. This paper proposes an effective way to build a graphical view of codes to better inform LLMs for code generation. To take the challenging structural graphical knowledge into LLMs, a meta-graph prompt is proposed for tuning-free models and a soft-prompting technique is proposed to inject the structural programming domain knowledge into the parameters of LLMs. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate code with improved accuracy and fluency.
Limitations
In this paper, we propose a graphical retrieval augmented generation method that can offer enhanced code generation. Despite the efficiency and effectiveness, there are also limitations within this work. For example, dependency on the quality of the external knowledge base could be a potential concern. The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes.
@@ Acknowledgments
This document has been adapted by Emily Allaway from the instructions for earlier ACL and NAACL proceedings, including those for NAACL 2024 by Steven Bethard, Ryan Cotterell and Rui Yan, ACL 2019 by Douwe Kiela and Ivan Vulić, NAACL 2019 by Stephanie Lukin and Alla Roskovskaya, ACL 2018 by Shay Cohen, Kevin Gimpel, and Wei Lu, NAACL 2018 by Margaret Mitchell and Stephanie Lukin, Bib suggestions for (NA)ACL 2017/2018 from Jason Eisner, ACL 2017 by Dan Gildea and Min-Yen Kan, NAACL 2017 by Margaret Mitchell, ACL 2012 by Maggie Li and Michael White, ACL 2010 by Jing-Shin Chang and Philipp Koehn, ACL 2008 by Johanna D. Moore, Simone Teufel, James Allan, and Sadaoki Furui, ACL 2005 by Hwee Tou Ng and Kemal Oflazer, ACL 2002 by Eugene Charniak and Dekang Lin, and earlier ACL and EACL formats written by several people, including John Chen, Henry S. Thompson and Donald Walker. Additional elements were taken from the formatting instructions of the International Joint Conference on Artificial Intelligence and the Conference on Computer Vision and Pattern Recognition.
Appendix A Implementation Details
Appendix B Prompt Template
Figure 4: Hard meta-graph prompt.
Figure 5: Soft prompting.